OpenStudy (compassionate):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). @mathmale

OpenStudy (the_fizicx99):

f(g(x)) = (x^2 - 7x - 9) - 2 Do the rest bby

OpenStudy (anonymous):

well let's see, since its f(g(-1)), its hould be set up like ....lol nvm tHe_FiZiCx99 got it

OpenStudy (anonymous):

if you want i could walk you through the rest

OpenStudy (compassionate):

Why did my x obliterate itself off Earth.

OpenStudy (anonymous):

basically plug in -1 in the g(x) equation. take the answer that u get and plug it into the f(x)

OpenStudy (compassionate):

Okay, so -3.

OpenStudy (compassionate):

f(x) = x - 2 f(x) = -3 - 2 = -5

OpenStudy (anonymous):

well from the looks of it you need to factor

OpenStudy (mathmale):

f(g(x)) = (x^2 - 7x - 9) - 2 was tHe_FiZiCx99's contribution and is correct. You could either plug in x = -1 now or simplify his expression first and then plug in -1.

OpenStudy (compassionate):

" (x^2 - 7x - 9) - 2" Okay but f(x) = x - 2 So, if we have " (x^2 - 7x - 9) - 2" Where did the x go?

OpenStudy (mathmale):

Just to be sure you're aware of this: f(g(x)) is a "composite function:" f(g(x)) = (x^2 - 7x - 9) - 2 The idea here is that you write this GENERAL composition. Once you have this, you replace every instance of x with the given value -1.

OpenStudy (mathmale):

f(g(x)) = (x^2 - 7x - 9) - 2 could be re-written as f(g(x)) = x^2 - 7x - 11. substituting -1 for x should produce the same final result as before.

OpenStudy (mathmale):

Some hints: (-1)^2 = 1 7(-1) = -7

OpenStudy (mathmale):

You asked: "" (x^2 - 7x - 9) - 2" Okay but f(x) = x - 2 So, if we have " (x^2 - 7x - 9) - 2" Where did the x go? " Answer: In f(x) = x - 2, you replace 'x' on both sides of the equation with g(x), which is x^2 - 7x - 9. Thus, f(x) = x - 2 becomes f(g(x)) = [x^2 - 7x - 9] - 2, and this reduces to f(g(x)) = x^2 - 7x - 11. Lastly, replace every instance of x by (-1).

OpenStudy (compassionate):

This is all so confusing. Could you explain it from stage one, like Functions 101. I know what you're saying but I don't understand why.

OpenStudy (compassionate):

"f(g(x)) = (x^2 - 7x - 9) - 2 could be re-written as f(g(x)) = x^2 - 7x - 11. " How can it be rewritten as -11 at the end? Where did that come from? Did we take it out of the parenthesis? Also, ho would we even solve it once we have f(g(x)) = x^2 - 7x - 11 And I'm still lost as to what happened to my lil x.

OpenStudy (anonymous):

Omg Comp ♥ lmfaooooo

OpenStudy (anonymous):

@compassionate so did you plug in -1 for g(x)?

OpenStudy (compassionate):

No. What do you mean?

OpenStudy (shamil98):

Stop trolling, moderators have more important things to do besides appease you. -.-

OpenStudy (the_fizicx99):

You tell 'em Sham <3

OpenStudy (anonymous):

so basically the question is asking you to find what the output is when you plug -1 in for g(x) and take that output and plug that in for x in f(x)

OpenStudy (compassionate):

Righto, but I have f(g(x)) = x^2 - 7x - 11 How would I even solve that?

OpenStudy (dan815):

hey comp can you post some nice math riddles after this question

OpenStudy (anonymous):

okay do my way first it will make sense to you.

OpenStudy (compassionate):

I don't understand your first way.

OpenStudy (anonymous):

hmm okay so plug in -1 in for g(x) what do u get?

OpenStudy (compassionate):

g(x) = x2 - 7x - 9 g(x) = -1 + 7 - 9 g(x) = -3

OpenStudy (anonymous):

is that x^2 if so a negative * negative is a positive right? so do u think ur answer is right?

OpenStudy (compassionate):

-2

OpenStudy (compassionate):

Yeah, I thought it would be positive, but my calculator said -1. Such lies. I feel attacked.

OpenStudy (anonymous):

show me how u got there

OpenStudy (anonymous):

haha yea for your calc u need to make sure u have parentheses.

OpenStudy (compassionate):

x2 - 7x - 9 -1^2 -7(-1) - 9 1 + 7 - 9 -1 DERP!

OpenStudy (anonymous):

good job! so now plug in -1 in f(x) and tell me what u get

OpenStudy (compassionate):

-3

OpenStudy (anonymous):

show me ur work :)

OpenStudy (compassionate):

f(x) = x - 2 So... -1 - 2

OpenStudy (anonymous):

yup! good job! thats ur asnwer:)

OpenStudy (compassionate):

That would I originally concluded, but I redid it and got 21 so I wasn't sure. I appreciate your time, tejasvir. You'd make a great Ambassador.

OpenStudy (anonymous):

np:) haha thanks:P